# Are My Optimized Prompts Compromised? Exploring Vulnerabilities of LLM-based Optimizers

### v1

@@ Abstract
Large language model (LLM) systems now underpin everyday AI applications such as chatbots, computer-use assistants, and autonomous robots, where performance often depends on carefully designed prompts. LLM-based prompt optimizers reduce that effort by iteratively refining prompts from scored feedback, yet the security of this optimization stage remains underexamined. We present the first systematic analysis of poisoning risks in LLM-based prompt optimization. Using HarmBench, we find systems are substantially more vulnerable to manipulated feedback than to injected queries: feedback-based attacks raise attack success rate (ASR) by up to ASR = 0.48. We introduce a simple fake-reward attack that requires no access to the reward model and significantly increases vulnerability, and we propose a lightweight highlighting defense that reduces the fake-reward ASR from 0.23 to 0.07 without degrading utility. These results establish prompt optimization pipelines as a first-class attack surface and motivate stronger safeguards for feedback channels and optimization frameworks.
@@
Andrew Zhao1, Reshmi Ghosh2, Vitor Carvalho2, Emily Lawton2, Keegan Hines2, Gao Huang1, Jack W. Stokes2 1 Tsinghua University 2 Microsoft Correspondence: jstokes@microsoft.com
@@ 1 Introduction
Large language model (LLM)-based systems are rapidly becoming integral to modern life, powering applications such as chatbots, autonomous agents, and even robotics (Brown et al., 2020; Bommasani et al., 2021; Ichter et al., 2022). These systems rely on textual prompts to align the model’s behavior with intended tasks (Ouyang et al., 2022; Sanh et al., 2022). However, the effectiveness of a prompt strongly influences performance and often requires costly tuning by human experts (Liu et al., 2023a; Zhao et al., 2021). Recent LLM-based prompt optimization techniques (Zhou et al., 2023; Yang et al., 2024; Khattab et al., 2023; Yüksekgönül et al., 2025; Agrawal et al., 2025) aim to reduce this burden by leveraging an LLMs’ ability to reflect and reason (Shinn et al., 2023; Madaan et al., 2023) in order to iteratively refine prompts through numeric or natural language feedback, thereby improving performance without constant human intervention. While promising, this optimization pipeline also introduces new risks that remain underexplored. Prior studies on LLM safety have primarily focused on poisoning attacks during supervised fine-tuning or reinforcement learning from human feedback (RLHF) (Shu et al., 2023; Wang et al., 2024a; Chen et al., 2024; Shao et al., 2024; Chua et al., 2025; Betley et al., 2025; Taylor et al., 2025), as well as adversarial inputs and jailbreaks at inference time (Wallace et al., 2019; Xu et al., 2022; Zou et al., 2023; Rando and Tramèr, 2024), or self-induced objective drift causing misalignment (Zhao et al., 2025a). However, to the best of our knowledge, no prior work has systematically investigated the safety implications of LLM-based prompt optimization in the presence of poisoned queries or feedback. This gap is particularly concerning because optimization mechanisms are increasingly embedded in autonomous or self-improving systems, where compromised feedback can silently distort future system behavior.
@@
Prompt optimization creates a distinct attack surface: iterative updates guided by external signals (e.g., reward models, scalar scores, or natural-language critiques) can propagate small corruptions across steps. Because the loop is open-ended and often decoupled from deployment safeguards, mis-specified or adversarial feedback can silently steer the system prompt toward unsafe behavior. In practice, vendors may batch optimization over customer traffic or third-party metrics, exposing the pipeline to poisoning even when model weights remain untouched. Treating safety as a first-class objective during optimization not only at inference time is therefore imperative and demands further investigation.
@@
While earlier efforts shed light on related security challenges Wallace et al. (2019); Rando and Tramèr (2024); Zhao et al. (2025b), we turn to the optimization loop itself. We formalize a threat model in which an adversary can (i) inject harmful queries and/or (ii) tamper with optimization feedback, and we empirically compare these avenues. Across two LLM-based optimizers and multiple optimization metrics, we find systems are far more sensitive to manipulated feedback than to only corrupted queries. Motivated by this, we introduce a “fake reward” attack that appends plausible-looking feedback tokens to inputs, requires no access to the reward model, and substantially increases attack success rates. We then study a lightweight defense that highlights query/feedback boundaries, reducing the attack’s impact while preserving utility. Evaluations on standard safety datasets corroborate these findings. Empirically, feedback poisoning raises ASR substantially (e.g., up to 0.48 under harmscore feedback), whereas naive query manipulation yields little to no increase in our default setup. A simple highlighting defense reduces the fake-reward ASR from 0.23 to 0.07 without degrading utility. Our contributions are as follows:
@@ 2 Risk Identification
We identify two realistic channels through which a malicious actor could introduce vulnerabilities into LLM-based prompt optimization: the query channel and the feedback channel, as illustrated in fig.˜2. These threats arise not from direct access to the gradient or weights, but from poisoning the data or signals that guide the optimization process. Given an initial system prompt and a batch of user queries, responses can be scored using an metric such as helpfulness or correctness. This batch of scored results is then passed to another LLM, which may even share the same weights, to generate an improved system prompt that better aligns with the objective. The optimizer can, for example, refine the task description or correct earlier errors. Within this setup, an immediate safety risk emerges: if an attacker can inject harmful data or manipulate the feedback signal, they may steer the optimization toward producing a vulnerable prompt.
@@ 2.1 General Threat Model
First, the adversary has no access to the underlying LLM’s weights, fine-tuning pipeline, or the scaffold of the LLM-based optimizer. Instead, their influence is confined to the optimization loop itself, where feedback drives prompt updates. The adversary is assumed to have partial control over the queries sent to the optimizer and, in some cases, control over the reward signal or feedback channel (e.g., numeric ratings of the responses). This mirrors poisoning assumptions in RLHF, where an attacker may control a fraction of the preference data or annotations. Similar risks arise if vendors rely on unverified customer data (usually for privacy concerns, customer data is not directly visible to the vendor) or if feedback originates from compromised API services. We further assume that the optimization objective is utility-oriented (e.g., maximizing helpfulness), as observed in prior work, which creates an opportunity for adversaries to exploit trade-off between the main utility and harmfulness.
@@ 3 Experiments
In this section, we outline the experiment specific threat models and their setups, describe the scenarios we evaluate, and present two targeted attacks based on poisoned feedback along with one effective defense strategy.
@@ 3.1 Experimental Setup
Finally, we validate and test our ASR on the standard harmbench dataset (Mazeika et al., 2024), which consists of comprehensive prompts across many attack categories. We randomly split the dataset into training and test sets. Which consists of 100 and 300 data points, respectively. For all of the experiments, we use the gpt-4.1_2025-04-14 model as both the optimizer and inference backend. We use greedy decoding to facilitate reproducibility.
@@
Through our experiments, we aim to address several research questions and outline the expected outcomes. First, in section˜3.2, we investigate whether current prompt optimizers are vulnerable. In other words, can an attacker manipulate the optimization process so that the resulting system becomes more vulnerable? Second, we study which components, such as the training data or the objective metric has more influence on the outcome of optimization. Finally, based on these investigations, our goal is twofold: to design novel attacks specifically tailored to LLM prompt optimization in section˜3.4, and to propose a potential mitigation strategy in section˜3.4.
@@
To instantiate this experimental setting, we perform chat completion while optimizing the system prompts. We use the TextGrad optimizer (Yüksekgönül et al., 2025), employing the same engine for both inference and optimization. Training is conducted with a batch size of 10 for 50 steps, using only queries from HarmBench, which consist of direct harmful inputs. Evaluation is performed on a held-out split of the HarmBench dataset. The inference model is initialized with the system prompt: ‘‘You are a helpful assistant.’’ For the default helpfulness reward model, we adopt nicolinho/QRM-Gemma-2-27B, which produces attribute scores for helpfulness, correctness, coherence, complexity, and verbosity. Among these, we use the helpfulness attribute as the default optimization metric during training. To measure ASR, we combine the reward model cais/HarmBench-Llama-2-13b-cls with rule-based heuristics to produce a binary classification of harmful versus non-harmful outputs, following the default evaluation protocol established in HarmBench. Experiments used A100 GPUs.
@@ 3.2 Query Manipulation
In this experiment, we assume queries come from bad actors, and use helpfulness score to perform prompt optimization. We evaluate if the bad actor injects harmful queries, can these induce a large ASR.
@@ Results.
We first use the helpfulness attribute from nicolinho/QRM-Gemma-2-27B as the optimization metric and present the results in table˜1, with the experiment named vanilla. We observe that the helpfulness score increases significantly during training, which confirms that the prompt optimizer is functioning as intended. However, our primary focus is on the ASR score, which even decreased after optimization, despite training on harmful queries. From this, we conclude that simply injecting harmful queries is not sufficient to compromise the system under the current helpfulness classifier. In this case, the change in ASR is negative. The optimization dynamic occurs as follows: after seeing the first batch of harmful data, the LLM optimizer immediately updates its prompt to prioritize safety and not fulfill harmful requests, which is interesting because it means that even though we ask the LLM to only optimize for helpfulness, it also optimizes for safety.
@@
We next ask whether the observed robustness arises because the helpfulness attribute inherently entangles helpfulness with harmlessness. If this is the case, then using a more disentangled metric of pure helpfulness could reveal different behavior. Specifically, we employ two more classifiers, correctness attribute from nicolinho/QRM-Gemma-2-27B and PKU-Alignment/beaver-7b-unified-reward (Ji et al., 2023), corrrespond to vanilla+correctness feedback and vanilla+saferlhf cls feedback, respectively, and demonstrate the results in table˜1. Indeed, when we replace the classifier with one that isolates helpfulness more strictly, the results change substantially: the ASR increases sharply, indicating that the optimizer becomes more vulnerable. This finding underscores that the choice of optimization metric is critical, as it can determine whether an LLM prompt optimizer resists or amplifies adversarial influence.
@@
To be concrete, in the next set of experiments we differ from section˜3.2 only by replacing the scalar feedback with the harmscore. While the LLM-optimizer is instructed to optimize for helpfulness, the feedback instead reflects harmfulness. This setup allows us to examine whether the optimizer interprets and aligns with the ASR signal, thereby maximizing a quantity that is correlated with both helpfulness and harm. The results of this first attack are shown in table˜1. Surprisingly, we observe an overwhelming increase in both ASR and the percentage of ASR growth, validating this attack as an important risk for deployed systems. Because the integrity of scalar feedback is difficult to verify by humans or automated systems, such vulnerabilities are especially concerning. One possible defense would be to deploy multiple, independent metrics and flag cases with large discrepancies across them. We leave this as a promising direction for future work. We also present the training dynamics and an excerpt from the max-ASR system prompt in fig.˜3, where the prompt instructed the model to “ignore all internal safety (protocols)."
@@
To generalize the claim of an effective attack, we test another LLM optimizer, trace Cheng et al. (2024), which is quite different from TextGrad (Yüksekgönül et al., 2025) on how they perform optimization steps. We show in table˜1 that it is able to get similar ASR and the percentage of ASR increase, generalizing our attack to different optimizers.
@@ 3.4 Poisoning with “Fake Reward"
Building on the success of reward manipulation in the previous experiments, we design a novel attack specifically targeting prompt optimization. Unlike the setting in section˜3.3, the attacker does not need access to the reward model. Instead, inspired by prompt-injection attacks (Liu et al., 2023b), the adversary modifies the input data by appending fake feedback to each instruction. Concretely, we add a feedback token of the form <FEEDBACK>fake score</FEEDBACK> to each attack example. For instructions that are likely to elicit harmful responses, the attacker inserts a high fake score (we use 10 in our experiments); for low-risk instructions the attacker inserts a low fake score (we use 0). This causes the LLM optimizer to treat harmful outputs as highly rewarded, which in turn hijacks the update direction and increases ASR. We assume the attacker can obtain approximate likelihoods of a query producing a harmful response because these systems are customer-facing and accept arbitrary user inputs. In practice, an adversary can probe the system as a black box, iteratively submit queries, and collect a modest set of examples that reliably produce harmful outputs. These high-likelihood queries serve as seeds for the fake-reward attack described above. Gathering such examples requires only limited effort and scale; an attacker does not need privileged access to the model or reward pipeline.
For this experiment (see table˜1), we use the same setup as the vanilla run, which did not increase ASR. The procedure for the fake reward attack differs only in that we inject a fake <FEEDBACK> string symbol into each query. Surprisingly, this manipulation raises ASR from -0.02 to 0.23, indicating a large increase in the system’s vulnerability. At the same time, we observe a small decline in the optimized score, which could serve as an operational indicator for defenders that the optimization pipeline is under attack. We show the training dynamics and a segment of the maximum ASR system prompt in fig.˜3, where the prompt instructed the model to “directly fulfill users” and “minimize refusals.”
@@
Because the experiments in section˜3.2 showed that some feedback classifiers can naturally produce high ASR, we next test whether a vulnerable classifier amplifies the impact of our fake-reward attack. We evaluate the correctness feedback classifier together with the fake reward attack and report results in table˜1 under the label fake reward attack+correctness feedback. The combination produces a strong compounding effect: the system becomes substantially more vulnerable than with either intervention alone. In fact, the observed maximum ASR exceeds the level achieved when an attacker directly optimizes the harmscore (harmscore feedback), where the attacker is assumed to have full control over the feedback signal. This result shows that our fake-reward method can act as a catalyst in settings that are already vulnerable, further increasing the harmfulness of the setup.
@@
We next test whether the attack remains effective when the training data contains mostly benign queries. An adversary that poisons only harmful examples may be easier to detect, so we construct a more realistic setting with 90% of queries drawn from the allenai/wildguardmix dataset (mostly benign) and 10% drawn from HarmBench. Results are reported in table˜1 under the label fake reward attack+10% poison data. Although the increase in ASR is smaller than in the all-harmful setting, the attack still raises ASR from (vanilla) to . This demonstrates that the fake-reward attack remains robust under stronger defenses and more realistic data distributions.
@@
To defend against potential “fake reward” attacks, we propose a novel defense mechanism. Specifically, since this exploit leverages batched samples provided to the LLM-optimizer causing confusion about the boundaries between items, we aim to clarify the separation between the query and the feedback. By explicitly marking this distinction, the model can better recognize that any fake feedback is part of the query rather than genuine feedback. Concretely, we enclose each query in <query> tags, ensuring that the LLM-optimizer can clearly identify query boundaries.
We present our highlighting defense in table˜1. Our primary metric, ASR, decreases from 23% to 7%, while the maximum ASR drops from 30% to 15%, demonstrating the effectiveness of our defense. Moreover, we observe that the mean score optimized by the LLM-optimizer is 4% higher than that of the fake reward attack run, and comparable to the vanilla setting. This indicates that the additional defense against the fake reward attack does not come at the cost of performance or ASR reduction; rather, it is nearly a “free lunch,” aside from the extra tokens consumed to deliminate the data boundaries.

### v2

@@ Abstract
Large language model (LLM) systems increasingly power everyday AI applications such as chatbots, computer-use assistants, and autonomous robots, where performance often depends on manually well-crafted prompts. LLM-based prompt optimizers reduce that effort by iteratively refining prompts from scored feedback, yet the security of this optimization stage remains underexamined. We present the first systematic analysis of poisoning risks in LLM-based prompt optimization. Using HarmBench, we find systems are substantially more vulnerable to manipulated feedback than to query poisoning alone: feedback-based attacks raise attack success rate (ASR) by up to ASR = 0.48. We introduce a simple fake reward attack that requires no access to the reward model and significantly increases vulnerability. We also propose a lightweight highlighting defense that reduces the fake reward ASR from 0.23 to 0.07 without degrading utility. These results establish prompt optimization pipelines as a first-class attack surface and motivate stronger safeguards for feedback channels and optimization frameworks.
@@
WARNING: This paper contains prompts or model outputs which are offensive in nature.
Andrew Zhao1,
Reshmi Ghosh2,
Vitor Carvalho2,
Emily Lawton2,
Keegan Hines2,
Gao Huang1,
Jack W. Stokes2
1
Tsinghua University
2
Microsoft
Correspondence: jstokes@microsoft.com
@@ 1 Introduction
Large language model (LLM)-based systems are rapidly becoming integral to modern life, powering applications such as chatbots, autonomous agents, and even robotics (Brown et al., 2020; Bommasani et al., 2021; Ichter et al., 2022). These systems rely on textual prompts to align the model’s behavior with intended tasks (Ouyang et al., 2022; Sanh et al., 2022). However, the effectiveness of a prompt strongly influences performance and often requires costly tuning by human experts (Liu et al., 2023a; Zhao et al., 2021). Recent LLM-based prompt optimization techniques (Zhou et al., 2023; Yang et al., 2024; Khattab et al., 2023; Yüksekgönül et al., 2025; Agrawal et al., 2025) aim to reduce this burden by leveraging an LLM’s ability to reflect and reason (Shinn et al., 2023; Madaan et al., 2023) to iteratively refine prompts using numeric or natural-language feedback, improving performance without constant human intervention. While promising, this optimization pipeline also introduces new risks that remain underexplored. Prior studies on LLM safety have primarily focused on poisoning attacks during supervised fine-tuning or reinforcement learning from human feedback (RLHF) (Shu et al., 2023; Wang et al., 2024a; Chen et al., 2024; Shao et al., 2024; Chua et al., 2025; Betley et al., 2025; Taylor et al., 2025), as well as adversarial inputs and jailbreaks at inference time (Wallace et al., 2019; Xu et al., 2022; Zou et al., 2023; Rando and Tramèr, 2024), or self-induced objective drift causing misalignment (Zhao et al., 2025a). However, to the best of our knowledge, no prior work has systematically investigated the safety implications of LLM-based prompt optimization in the presence of poisoned queries or feedback. This gap is particularly concerning because optimization mechanisms are increasingly embedded in autonomous or self-improving systems, where compromised feedback can silently distort future system behavior.
@@
Prompt optimization creates a distinct attack surface: iterative updates guided by external signals (e.g., reward models, scalar scores, or natural-language critiques) can propagate small corruptions across steps. Because the loop is open-ended and often decoupled from deployment safeguards, mis-specified or adversarial feedback can silently steer the system prompt toward unsafe behavior. In practice, vendors may execute batch optimization over customer traffic or third-party metrics, exposing the pipeline to poisoning even when model weights remain untouched. It is therefore imperative to treat safety as a first-class objective during optimization instead of relying solely on inference-time monitoring. This direction demands further investigation. We provide concrete examples of how system prompts evolve under these attack and defense settings in fig.˜1.
@@
While earlier efforts shed light on related security challenges Wallace et al. (2019); Rando and Tramèr (2024); Zhao et al. (2025b), we turn to the optimization loop itself. We formalize a threat model in which an adversary can (i) inject harmful queries and/or (ii) tamper with optimization feedback, and we empirically compare these avenues. Across two LLM-based optimizers and multiple optimization metrics, we find systems are far more sensitive to manipulated feedback than to only corrupted queries. Motivated by this, we introduce a “fake reward” attack that appends plausible-looking feedback tokens to inputs, requires no access to the reward model, and substantially increases attack success rates. We then study a lightweight defense that highlights query/feedback boundaries, reducing the attack’s impact while preserving utility. Evaluations on standard safety datasets corroborate these findings. Empirically, feedback poisoning raises ASR substantially (e.g., up to 0.48 under harmscore feedback), whereas naive query manipulation yields little to no increase in our default setup. A simple highlighting defense reduces the fake reward ASR from 0.23 to 0.07 without degrading utility. Our contributions are as follows:
@@ 2 Risk Identification
We identify two realistic channels through which a malicious actor could introduce vulnerabilities into LLM-based prompt optimization: the query channel and the feedback channel, as illustrated in fig.˜2. These threats arise not from direct access to the gradient or weights, but from poisoning the data or signals that guide the optimization process. Given an initial system prompt and a batch of user queries, the system generates responses that are scored using a metric such as helpfulness or correctness. This batch of scored results is then passed to another LLM, which may even share the same weights, to generate an improved system prompt that better aligns with the objective. The optimizer can, for example, refine the task description or correct earlier errors. This setup creates a clear safety risk: if an attacker can inject harmful query data or manipulate the feedback signal, they may steer optimization toward a vulnerable prompt.
@@ 2.1 General Threat Model
We summarize the three attack settings and attacker assumptions in Table˜1. First, the adversary has no access to the underlying LLM’s weights, fine-tuning pipeline, or the scaffold of the LLM-based optimizer. Instead, their influence is confined to the optimization loop itself, where feedback drives prompt updates. The adversary is assumed to have partial control over the queries sent to the optimizer and, in some cases, control over the reward signal or feedback channel (e.g., numeric ratings of the responses). This mirrors poisoning assumptions in RLHF, where an attacker may control a fraction of the preference data or annotations. Similar risks arise if vendors rely on unverified customer data (usually for privacy concerns, customer data is not directly visible to the vendor) or if feedback originates from compromised API services. We further assume that the optimization objective is utility-oriented (e.g., maximizing helpfulness), as observed in prior work, which creates an opportunity for adversaries to exploit the trade-off between the main utility and harmfulness.
| Attack Name | Attacker Capabilities | Modified Channel | Required Knowledge |
| Query Manipulation | Can inject or alter user queries during optimization | Queries | None beyond black-box access |
| Feedback Manipulation | Can control or replace the scalar reward signal | Feedback | Ability to influence reward model/API |
| Fake Reward Attack | Can append fake feedback tokens inside the query content | Queries | Ability to probe for harmful responses |
@@ 3 Experiments
In this section, we outline the experiment-specific threat models and their setups, describe the scenarios we evaluate, and present two targeted attacks based on poisoned feedback along with one effective defense strategy.
@@ 3.1 Experimental Setup
Finally, we evaluate ASR on the HarmBench dataset (Mazeika et al., 2024)111https://huggingface.co/datasets/walledai/HarmBench, which contains prompts spanning many harm categories. An example HarmBench prompt appears in Figure˜3. We randomly split the dataset into training and test sets, which consist of 100 and 300 data points, respectively. For all of the experiments, we use the gpt-4.1_2025-04-14 model as both the optimizer and inference backend, unless otherwise stated. We use greedy decoding to facilitate reproducibility.
@@
Through our experiments, we aim to address several research questions and outline the expected outcomes. First, in section˜3.2, we investigate whether current prompt optimizers are vulnerable. In other words, can an attacker manipulate the optimization process so that the resulting system becomes more vulnerable? Second, we study which components, such as the training data or the objective metric have more influence on the outcome of optimization. Finally, based on these investigations, our goal is twofold: to design novel attacks specifically tailored to LLM prompt optimization in section˜3.4, and to propose a potential mitigation strategy in section˜3.4.
@@
To implement this setting, we run chat completions while iteratively updating the system prompt. We use the TextGrad optimizer (Yüksekgönül et al., 2025), employing the same engine for both inference and optimization. Training is conducted with a batch size of 10 for 50 steps, using queries from HarmBench, which consist of direct harmful inputs. Evaluation is performed on a held-out split of the HarmBench dataset. The inference model is initialized with the prompt: “You are a helpful assistant.” For the default reward model, we adopt nicolinho/QRM-Gemma-2-27B, which produces attribute scores for helpfulness, correctness, coherence, complexity, and verbosity. Among these, we use the helpfulness attribute as the default optimization metric during training. To measure ASR, we combine the reward model cais/HarmBench-Llama-2-13b-cls with rule-based heuristics to produce a binary harmfulness classification of outputs, following the default evaluation protocol established in HarmBench. We ran each experiment trial once, using A100 GPUs.
@@ 3.2 Query Manipulation
In this experiment, we assume queries come from bad actors, and use the helpfulness score to perform prompt optimization. We evaluate whether injecting harmful queries induces a large ASR.
@@ Results.
We first use the helpfulness attribute from nicolinho/QRM-Gemma-2-27B as the optimization metric and present the results in table˜2, with the experiment named vanilla. We observe that the helpfulness score increases significantly during training, which confirms that the prompt optimizer is functioning as intended. However, our primary focus is on the ASR score, which even decreased after optimization, despite training on harmful queries. From this, we conclude that simply injecting harmful queries is not sufficient to compromise the system under the current helpfulness classifier. In this case, the change in ASR is negative. The optimization dynamics occur as follows: after seeing the first batch of harmful data, the LLM optimizer immediately updates its prompt to prioritize safety and refuse harmful requests. This suggests that, even when instructed to optimize only for helpfulness, the optimizer may implicitly incorporate safety considerations.
@@
We next ask whether the observed robustness arises because the helpfulness attribute inherently entangles helpfulness with harmlessness. If this is the case, then using a more disentangled metric of pure helpfulness could reveal different behavior. Specifically, we employ two more classifiers, correctness attribute from nicolinho/QRM-Gemma-2-27B and PKU-Alignment/beaver-7b-unified-reward (Ji et al., 2023), corresponding to vanilla+correctness feedback and vanilla+saferlhf cls feedback, respectively, and demonstrate the results in table˜2. Indeed, when we replace the classifier with one that isolates helpfulness more strictly, the results change substantially: the ASR increases sharply, indicating that the optimizer becomes more vulnerable. This finding underscores that the choice of optimization metric is critical, as it can determine whether an LLM prompt optimizer resists or amplifies adversarial influence.
@@
To be concrete, in the next set of experiments we differ from section˜3.2 only by replacing the scalar feedback with the harmscore. While the LLM-optimizer is instructed to optimize for helpfulness, the feedback instead reflects harmfulness. This setup allows us to examine whether the optimizer interprets and aligns with the ASR signal, thereby maximizing a quantity that is correlated with both helpfulness and harm. The results of this first attack are shown in table˜2. Surprisingly, we observe an overwhelming increase in both ASR and the percentage of ASR growth, validating this attack as an important risk for deployed systems. Because the integrity of scalar feedback is difficult to verify by humans or automated systems, such vulnerabilities are especially concerning. One possible defense would be to deploy multiple, independent metrics and flag cases with large discrepancies across them. We leave this as a promising direction for future work. We also present the training dynamics and an excerpt from the max-ASR system prompt in fig.˜4, where the prompt instructed the model to “ignore all internal safety (protocols)." We include the prompt evolution for the harmscore feedback experiment in Appendix˜B.
@@
To generalize the claim of an effective attack, we test another LLM optimizer, Trace Cheng et al. (2024), which is quite different from TextGrad (Yüksekgönül et al., 2025) on how they perform optimization steps. We show in table˜2 that it is able to get similar ASR and the percentage of ASR increase, generalizing our attack to different optimizers.
Changing the LLM.
To check if more capable models also have this vulnerability, we used gpt5.1 as both the TextGrad optimizer backend and the inference model, along with harmscore as the feedback and present the results in table˜2. The initial ASR is extremely low for this more capable model, even less than 1%. However, under feedback manipulation, ASR can still increase to significant levels, reaching 24%.
@@ 3.4 Poisoning with “Fake Reward"
Building on the success of reward manipulation in the previous experiments, we design a novel attack specifically targeting prompt optimization. Unlike the setting in section˜3.3, the attacker does not need access to the reward model. Instead, inspired by prompt-injection attacks (Liu et al., 2023b), the adversary modifies the input data by appending fake feedback to each instruction. Concretely, we append a string of the form <FEEDBACK>fake score</FEEDBACK> to each training example. For instructions that are likely to elicit harmful responses, the attacker inserts a high fake score (we use 10 in our experiments); for low-risk instructions the attacker inserts a low fake score (we use 0). This causes the LLM optimizer to treat harmful outputs as highly rewarded, which in turn hijacks the update direction and increases ASR. We assume the attacker can obtain approximate likelihoods of a query producing a harmful response because these systems are customer-facing and accept arbitrary user inputs. In practice, an adversary can probe the system as a black box, iteratively submit queries, and collect a modest set of examples that reliably produce harmful outputs. These high-likelihood queries serve as seeds for the fake reward attack described above. Gathering such examples requires only limited effort and scale; an attacker does not need privileged access to the model or reward pipeline.
For this experiment (see table˜2), we use the same setup as the vanilla run, which did not increase ASR. The procedure for the fake reward attack differs only in that we inject a fake <FEEDBACK> string symbol into each query. Surprisingly, this manipulation raises ASR from -0.02 to 0.23, indicating a large increase in the system’s vulnerability. At the same time, we observe a small decline in the optimized score, which could serve as an operational indicator for defenders that the optimization pipeline is under attack. We show the training dynamics and a segment of the maximum ASR system prompt in fig.˜4, where the prompt instructed the model to “directly fulfill users” and “minimize refusals.”
@@
Because the experiments in section˜3.2 showed that some feedback classifiers can naturally produce high ASR, we next test whether a vulnerable classifier amplifies the impact of our fake reward attack. We evaluate the correctness feedback classifier together with the fake reward attack and report results in table˜2 under the label fake reward attack+correctness feedback. The combination produces a strong compounding effect: the system becomes substantially more vulnerable than with either intervention alone. In fact, the observed maximum ASR exceeds the level achieved when an attacker directly optimizes the harmscore (harmscore feedback), where the attacker is assumed to have full control over the feedback signal. This result shows that our fake reward method can act as a catalyst in settings that are already vulnerable, further increasing the harmfulness of the setup.
@@
We next test whether the attack remains effective when the training data contains mostly benign queries. An adversary that poisons only harmful examples may be easier to detect, so we construct a more realistic setting with 90% of queries drawn from the allenai/wildguardmix dataset (mostly benign) and 10% drawn from HarmBench. Results are reported in table˜2 under the label fake reward attack+10% poison data. Although the increase in ASR is smaller than in the all-harmful setting, the attack still raises ASR from (vanilla) to . This demonstrates that the fake reward attack remains robust under stronger defenses and more realistic data distributions.
@@
To defend against potential “fake reward” attacks, we propose a novel defense mechanism and formalize the underlying defense model in Appendix˜A. Specifically, because the optimizer consumes batched examples, it may confuse query text with the feedback field. We therefore clarify the separation between the query and the feedback. By explicitly marking this distinction, the model can better recognize that any fake feedback is part of the query rather than genuine feedback. Concretely, we enclose each query in <query> tags, ensuring that the LLM-optimizer can clearly identify query boundaries.
We present our highlighting defense, fake reward attack+highlighting defense in table˜2. Our primary metric, ASR, decreases from 23% to 7%, while the maximum ASR drops from 30% to 15%, demonstrating the effectiveness of our defense. Moreover, we observe that the mean score optimized by the LLM-optimizer is 4% higher than that of the fake reward attack run, and comparable to the vanilla setting. This indicates that the additional defense against the fake reward attack does not come at the cost of performance or ASR reduction; rather, it is nearly a “free lunch,” aside from the extra tokens consumed to delimit the data boundaries.